# A generalization of the relaxation-based matrix splitting iterative method for solving the system of generalized absolute value equations

### v1

@@ 3 Convergence analysis
then the GAVE (1.1) has a unique solution $x^{*}$ and the sequence $\{x^{k}\}_{k=0}^{\infty}$ generated by the MSOR iteration (2.12) converges to $x^{*}$.
We can have the following result from Corollary 3.14 with $B=I$.
Corollary 3.15.
@@
$-\frac{2\omega_{\rm MSOR}|1-\alpha_{\rm MSOR}\omega_{\rm MSOR}|}{|2-\omega_{\rm MSOR}|}\|A^{-1}\|\|A\|\bigg{|}<1$ (3.65)
@@
$\qquad<\Big{(}1-|1-\alpha_{\rm MSOR}\omega_{\rm MSOR}|\Big{)}\Big{(}1-|1-\omega_{\rm MSOR}|-\omega_{\rm MSOR}|1-\frac{2}{2-\omega_{\rm MSOR}}|\Big{)},$ (3.66)
then the AVE (1.2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{k}\}_{k=0}^{\infty}$ generated by the MSOR iteration (2.4) converges to $x^{*}$.
Remark 3.10.
The conditions (3.65) and (3.66) are different from the conditions of [7, Theorem 3.2] and the unique existence of the solution of the AVEs (1.2) is not explored there.
@@
Corollary 3.16.
Let $A,B\in\mathbb{R}^{n\times n}$ be two given nonsingular matrices, $\theta=\omega_{\rm NSOR}>0,~{}\sigma_{\rm NSOR}>0$. If
$\Big{|}(1-\omega_{\rm NSOR})^{2}+\frac{|1-\omega_{\rm NSOR}|\omega_{\rm NSOR}}{\sigma_{\rm NSOR}}\Big{[}\|\sigma_{\rm NSOR}I-B\|-\|A\|\|A^{-1}B\|\Big{]}\Big{|}<1~{}$
@@
$\frac{\omega_{\rm NSOR}}{\sigma_{\rm NSOR}}\|A^{-1}B\|\Big{(}\omega_{\rm NSOR}\|B\|+2|1-\omega_{\rm NSOR}|\|A\|\Big{)}$
$\qquad<\Big{(}1-|1-\omega_{\rm NSOR}|\Big{)}\Big{(}1-|1-\omega_{\rm NSOR}|-\frac{\omega_{\rm NSOR}}{\sigma_{\rm NSOR}}\|\sigma_{\rm NSOR}I-B\|\Big{)},$
then the GAVE (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{k}\}_{k=0}^{\infty}$ generated by the NSOR iteration (2.13) converges to $x^{*}$.
When $B=I$ then the Corollary 3.16 turns into Corollary 3.17.
Corollary 3.17.
@@
$\Big{|}(1-\omega_{\rm NSOR})^{2}+\frac{|1-\omega_{\rm NSOR}|\omega_{\rm NSOR}}{\sigma_{\rm NSOR}}\Big{[}|1-\sigma_{\rm NSOR}|-\|A\|\|A^{-1}\|\Big{]}\Big{|}<1~{}$ (3.67)
@@
$\frac{\omega_{\rm NSOR}}{\sigma_{\rm NSOR}}\|A^{-1}\|\Big{(}\omega_{\rm NSOR}+2|1-\omega_{\rm NSOR}|\|A\|\Big{)}$
$\qquad<\Big{(}1-|1-\omega_{\rm NSOR}|\Big{)}\Big{(}1-|1-\omega_{\rm NSOR}|-\frac{\omega_{\rm NSOR}}{\sigma_{\rm NSOR}}|1-\sigma_{\rm NSOR}|\Big{)},$ (3.68)
then the AVEs (1.2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{k}\}_{k=0}^{\infty}$ generated by the NSOR iteration (1.9) converges to $x^{*}$.
Remark 3.11.
The convergence conditions (3.67) and (3.68) are new, which are not appeared in [5]. In addition, the existence and uniqueness of the solution to the AVEs (1.2) is not explored in [5].
For FPI-SS iteration [15], the following corollary is obtained from Corollary 3.1.
Corollary 3.18.
@@
[…]
@@
then the GAVEs (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the FPI-SS iteration (2.7) converges to $(x^{*},y^{*}=|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
Remark 3.12.
@@
In [15, Theorem 1], under the assumption that the GAVEs (1.1) has a unique solution, the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the FPI-SS iteration converges to the unique solution $x^{*}$ of the GAVEs (1.1) if
[…]
@@
It is easy to see that $\dfrac{2}{1+\hat{a}+\hat{b}}<\dfrac{2(1-\hat{a})}{1-\hat{a}+\hat{b}}$, which implies that (3.69) is weaker than (3.70). Furthermore we can conclude that the GAVEs (1.1) has a unique solution under condition (3.70).

### v2

@@ 3 Convergence analysis
then GAVE (1) has a unique solution $x^{*}$ and the sequence $\{x^{k}\}_{k=0}^{\infty}$ generated by the MSOR iteration (22) converges to $x^{*}$.
We can have the following result from Corollary 14 with $B=I$.
Corollary 15.
@@
$-\frac{2\omega_{\rm MSOR}|1-\alpha_{\rm MSOR}\omega_{\rm MSOR}|}{|2-\omega_{\rm MSOR}|}\|A^{-1}\|\|A\|\bigg{|}<1$ (71)
@@
$\qquad<\Big{(}1-|1-\alpha_{\rm MSOR}\omega_{\rm MSOR}|\Big{)}\Big{(}1-|1-\omega_{\rm MSOR}|-\omega_{\rm MSOR}|1-\frac{2}{2-\omega_{\rm MSOR}}|\Big{)},$ (72)
then AVE (2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{k}\}_{k=0}^{\infty}$ generated by the MSOR iteration (14) converges to $x^{*}$.
Remark 10.
The conditions (71) and (72) are different from the conditions of [8, Theorem 3.2] and the unique existence of the solution of AVEs (2) is not explored there.
@@
Corollary 16.
Let $A,B\in\mathbb{R}^{n\times n}$ be two given nonsingular matrices, $\omega_{\rm NSOR}>0,~{}\sigma_{\rm NSOR}>0$. If
$\Big{|}|1-\omega_{\rm NSOR}|\|I-\frac{\omega_{\rm NSOR}^{2}}{\sigma_{\rm NSOR}}B\|-\frac{\omega_{\rm NSOR}^{2}|1-\omega_{\rm NSOR}|}{\sigma_{\rm NSOR}}\|A^{-1}B\|\|A\|\Big{|}<1~{}$ (73a)
@@
$\frac{\omega_{\rm NSOR}^{2}}{\sigma_{\rm NSOR}}\|A^{-1}B\|\Big{(}\|B\|+|1-\omega_{\rm NSOR}|\|A\|\Big{)}<\Big{(}1-|1-\omega_{\rm NSOR}|\Big{)}\Big{(}1-\|I-\frac{\omega_{\rm NSOR}^{2}}{\sigma_{\rm NSOR}}B\|\Big{)},$ (73b)
then GAVE (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{k}\}_{k=0}^{\infty}$ generated by the NSOR iteration (23) converges to $x^{*}$.
When $B=I$, the Corollary 16 turns into Corollary 17.
Corollary 17.
@@
$\Big{|}|1-\omega_{\rm NSOR}||1-\frac{\omega_{\rm NSOR}^{2}}{\sigma_{\rm NSOR}}|-\frac{\omega_{\rm NSOR}^{2}|1-\omega_{\rm NSOR}|}{\sigma_{\rm NSOR}}\|A^{-1}\|\|A\|\Big{|}<1~{}$ (74a)
@@
$\frac{\omega_{\rm NSOR}^{2}}{\sigma_{\rm NSOR}}\|A^{-1}\|\Big{(}1+|1-\omega_{\rm NSOR}|\|A\|\Big{)}<\Big{(}1-|1-\omega_{\rm NSOR}|\Big{)}\Big{(}1-|1-\frac{\omega_{\rm NSOR}^{2}}{\sigma_{\rm NSOR}}|\Big{)},$ (74b)
then AVEs (2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{k}\}_{k=0}^{\infty}$ generated by the NSOR iteration (9) converges to $x^{*}$.
Remark 11.
The convergence conditions (74a) and (74b) are new, which are not appeared in [6]. In addition, the existence and uniqueness of the solution to AVEs (2) is not explored in [6].
For FPI-SS iteration [16], the following corollary is obtained from Corollary 1.
Corollary 18.
@@
$\hat{a}+\hat{b}<1\quad\text{and}\quad 0<\omega_{\rm FPI\text{-}SS}<\frac{2(1-\hat{a})}{1-\hat{a}+\hat{b}},$ (75)
@@
then GAVEs (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the FPI-SS iteration (17) converges to $(x^{*},y^{*}=|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
Remark 12.
@@
In [16, Theorem 1], under the assumption that GAVEs (1) has a unique solution, the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the FPI-SS iteration converges to the unique solution $x^{*}$ of GAVEs (1) if
$\hat{a}+\hat{b}<1\quad\text{and}\quad 0<\omega_{\rm FPI\text{-}SS}<\dfrac{2}{1+\hat{a}+\hat{b}}.$ (76)
@@
It is easy to see that $\dfrac{2}{1+\hat{a}+\hat{b}}<\dfrac{2(1-\hat{a})}{1-\hat{a}+\hat{b}}$, which implies that (75) is weaker than (76). Furthermore we can conclude that GAVEs (1) has a unique solution under condition (76).
4 Comparison theorems
It follows from (37) that the smaller value of $\rho(T)$ is (where $T$ is defined as in (32)), the faster the GRMS iteration (13) will converge later on. In this section, we will theoretically compare the convergence rate of the GRMS iteration (13) with some existing iterations, i.e., the RMS iteration (3), the MGSOR iteration (19) and the NSOR iteration (23). To this end, we introduce the following two lemmas.
Lemma 1 ([2, p. 27]).
For $n\geq 2$, a nonnegative matrix $U\in\mathbb{R}^{n\times n}$ is irreducible if $(I+U)^{n-1}$ is a positive matrix, i.e. $(I+U)^{n-1}>0$.
Lemma 2 ([2, p. 27]).
Let $U,R\in\mathbb{R}^{n\times n}$. If $R+U$ is an irreducible matrix, $U\neq R$ and $U\geq R\geq 0$, $\rho(U)>\rho(R)$.
4.1 GRMS vs. RMS
Let
$T_{\rm RMS}=\begin{bmatrix}\|M^{-1}_{\rm RMS}N_{\rm RMS}\|&\|M^{-1}_{\rm RMS}B\|\\ \tau_{\rm RMS}\|M^{-1}_{\rm RMS}N_{\rm RMS}\|&|1-\tau_{\rm RMS}|+\tau_{\rm RMS}\|M^{-1}_{\rm RMS}B\|\\ \end{bmatrix},$
which is derived from (32) with $M=M_{\rm RMS}$, $Q=I$, $Q_{1}=\frac{\theta}{\tau_{\rm RMS}}I$, $H=0$ and $\theta=\tau_{\rm RMS}$. Similar to the proof of Theorem 1, the smaller value of $\rho(T_{\rm RMS})$ is, the faster the RMS iteration (3) will converge later on. For GRMS and RMS, we have the following comparison theorem.
Theorem 2.
If $A,B\in\mathbb{R}^{n\times n}$ with $A=M-N$ such that $M$ is nonsingular and $M_{\rm RMS}=M$. $Q=Q_{1}-Q_{2}\in\mathbb{R}^{n\times n}$ and $Q_{1}$ is nonsingular. $\tau_{\rm RMS}>0$ and $\theta>0$. Assume that notation (25) and [27, Lemma 3] hold, and $T$ is an irreducible matrix. If
$\|Q\|\leq 1,~{}\beta c<|1-\tau_{\rm RMS}|+\tau_{\rm RMS}\|M^{-1}B\|+1,~{}\beta(a+1)<\tau_{\rm RMS}a,~{}\alpha c+d<1,$ (77a)
$(1-d)[\tau_{\rm RMS}a-\beta(a+1)]>\alpha c(\tau_{\rm RMS}a-\beta)+a\alpha[1-|1-\tau_{\rm RMS}|-\tau_{\rm RMS}\|M^{-1}B\|]$ (77b)
and
$\footnotesize\dfrac{1+\beta c-(|1-\tau_{\rm RMS}|+\tau_{\rm RMS}\|M^{-1}B\|)}{1-(\alpha c+d)}<\theta<\min\left\{\dfrac{|1-\tau_{\rm RMS}|+\tau_{\rm RMS}\|M^{-1}B\|+1-\beta c}{1+\alpha c+d},~{}\frac{\tau_{\rm RMS}a-(a+1)\beta}{a\alpha}\right\},$ (77c)
then $\rho(T)<\rho(T_{\rm RMS})<1$.
Proof.
According to Lemma 2, condition (77) and the assumption that [27, Lemma 3] hold, we know that $0\leq T\leq T_{\rm RMS}$ and $T\neq T_{\rm RMS}$, which implied that $\rho(T)<\rho(T_{\rm RMS})<1$ since $T$ is an irreducible matrix. ∎
The following example is used to verify Theorem 2.
Example 1.
Consider GAVEs (1) with $A=M-N$, $B=\begin{bmatrix}2&-1.5&1.5&0\\ -1.5&2&0&1.5\\ 1.5&0&2&-1.5\\ 0&1.5&-1.5&2\\ \end{bmatrix}$ and $b=Ax^{*}-B|x^{*}|$ where $x^{*}=[-\frac{1}{2},0,-\frac{1}{2},0]^{\top}$.
  1. [(i)]
  2. 1.
    For GRMS iteration: $M=\begin{bmatrix}8&-1.5&1.5&0\\ -1.5&8&0&1.5\\ 1.5&0&8&-1.5\\ 0&1.5&-1.5&8\\ \end{bmatrix}$, $N=-\dfrac{1}{8}I$, $Q=Q_{1}-Q_{2}=0.98I$ with $Q_{1}=0.97I$, $H=0.01I$ and $\theta=0.66$. Then $T\approx\begin{bmatrix}0.0250&0.4455\\ 0.0276&0.6545\end{bmatrix}$ is an irreducible matrix and $\rho(T)\approx 0.6734$;
  3. 2.
    For RMS iteration: $M_{\rm RMS}=M$, $N_{\rm RMS}=N$ and $\tau_{\rm RMS}=1.21$. Then $T_{\rm RMS}\approx\begin{bmatrix}0.0250&0.4545\\ 0.0302&0.7600\end{bmatrix}$ and $\rho(T_{\rm RMS})\approx 0.7783$.
It’s easy to verify that the conditions of Theorem 2 hold and $\rho(T)<\rho(T_{\rm RMS})<1$. By setting initial vector $x^{(0)}=y^{(0)}=[0,0,0,0]^{\top}$, we can see that the IT (the number of iteration) of GRMS is less than that of RMS from Figure 1.
Figure 1: The curve of RES:=$\frac{\|Ax^{(k)}-B|x^{(k)}|-b\|_{2}}{\|b\|_{2}}$ and IT (the number of iteration) between GRMS and RMS iteration. The methods are terminated when ${\rm RES}\leq 10^{-9}$.
4.2 GRMS vs. MGSOR
Let
$\footnotesize T_{\rm MGSOR}=\begin{bmatrix}|1-\alpha_{\rm MGSOR}|&\alpha_{\rm MGSOR}\|A^{-1}BQ_{\rm MGSOR}\|\\ \beta_{\rm MGSOR}|1-\alpha_{\rm MGSOR}|\|Q^{-1}_{\rm MGSOR}\|&\alpha_{\rm MGSOR}\beta_{\rm MGSOR}\|A^{-1}BQ_{\rm MGSOR}\|\|Q_{\rm MGSOR}^{-1}\|+|1-\beta_{\rm MGSOR}|\\ \end{bmatrix},$
which is derived from (32) with $M=\frac{1}{\alpha_{\rm MGSOR}}A$, $Q=Q_{\rm MGSOR},Q_{1}=\frac{\theta}{\beta_{\rm MGSOR}}Q_{\rm MGSOR}$, $H=0$ and $\theta=\beta_{\rm MGSOR}$. Similar to the proof of Theorem 1, the smaller value of $\rho(T_{\rm MGSOR})$ is, the faster the MGSOR iteration (15) will converge later on. For GRMS and MGSOR, we have the following comparison theorem.
Theorem 3.
Let $A,B\in\mathbb{R}^{n\times n}$ with $A=M-N$ such that $M$ is nonsingular. $Q_{\rm MGSOR}\in\mathbb{R}^{n\times n}$, $Q=Q_{1}-Q_{2}\in\mathbb{R}^{n\times n}$ and $Q_{1}$ is nonsingular. $\alpha_{\rm MGSOR},\beta_{\rm MGSOR}$ and $\theta>0$. Assume that notation (25) and condition (63) hold, and $T$ is an irreducible matrix.
If
$a<|1-\alpha_{\rm MGSOR}|,~{}c<\alpha_{\rm MGSOR}\|A^{-1}BQ_{\rm MGSOR}\|,~{}\alpha c+d<1$ (78a)
and
$\beta\big{(}(a+1)(1-d)-\alpha c\big{)}-\beta_{\rm MGSOR}|1-\alpha_{\rm MGSOR}|\|Q^{-1}_{\rm MGSOR}\|(1-\alpha c-d)<$ (78b)
$a\alpha\left[|1-\beta_{\rm MGSOR}|+\alpha_{\rm MGSOR}\beta_{\rm MGSOR}\|A^{-1}BQ_{\rm MGSOR}\|\|Q^{-1}_{\rm MGSOR}\|-1\right]$ (78c)
such that
$\tilde{\mu}<\theta<\min\left\{\tilde{\nu},~{}\Delta\right\},$ (78d)
where
$\tilde{\mu}=\frac{1+\beta c-\left(|1-\beta_{\rm MGSOR}|+\alpha_{\rm MGSOR}\beta_{\rm MGSOR}\|A^{-1}BQ_{\rm MGSOR}\|\|Q^{-1}_{\rm MGSOR}\|\right)}{1-(\alpha c+d)},$
$\tilde{\nu}=\frac{1-\beta c+|1-\beta_{\rm MGSOR}|+\alpha_{\rm MGSOR}\beta_{\rm MGSOR}\|A^{-1}BQ_{\rm MGSOR}\|\|Q^{-1}_{\rm MGSOR}\|}{1+\alpha c+d}$
and
$\Delta=\frac{\beta_{\rm MGSOR}|1-\alpha_{\rm MGSOR}|\|Q^{-1}_{\rm MGSOR}\|-(a+1)\beta}{a\alpha},$
then $\rho(T)<\rho(T_{\rm MGSOR})<1$.
Proof.
According to Lemma 2, condition (78) and the assumption that (63) hold, we know that $0\leq T<T_{\rm MGSOR}$, which implied that $\rho(T)<\rho(T_{\rm MGSOR})<1$ since $T$ is an irreducible matrix. ∎
The following Example 2 is used to verify Theorem 3.
Example 2.
Consider AVEs (2) with $A=M-N$ and $b=Ax^{*}-|x^{*}|$ with $x^{*}=[\frac{1}{2},0,\frac{1}{2},0]^{\top}$.
  1. [(i)]
  2. 1.
    For GRMS iteration: $M=\begin{bmatrix}24&-1&-1&0\\ -1&24&0&-1\\ -1&0&24&-1\\ 0&-1&-1&24\\ \end{bmatrix}$, $N=\dfrac{1}{4}I$, $Q=0.982I$, $Q_{1}=0.981I$, $H=0.01I$ and $\theta=1$. Then $T\approx\begin{bmatrix}0.0114&0.0446\\ 0.0219&0.0470\end{bmatrix}$ is an irreducible matrix and $\rho(T)\approx 0.0651$;
  3. 2.
    For MGSOR iteration: $Q_{\rm MGSOR}=Q$ and $\alpha_{\rm MGSOR}=1.05,~{}\beta_{\rm MGSOR}=1.07$. Then $T_{\rm MGSOR}\approx\begin{bmatrix}0.0500&0.0474\\ 0.0545&0.1217\end{bmatrix}$ and $\rho(T_{\rm MGSOR})\approx 0.1480$.
We can verify that the conditions of Theorem 3 are satisfied and we have $\rho(T)<\rho(T_{\rm MGSOR})<1$. Set initial vectors $x^{(0)}=y^{(0)}=[0,0,0,0]^{\top}$. Numerical results for Example 2 are reported in Figure 2, from which we can conclude that the number of iteration for the GRMS method is less than that of the MGSOR method.
Figure 2: The curve of RES:=$\frac{\|Ax^{(k)}-|x^{(k)}|-b\|_{2}}{\|b\|_{2}}$ for Example 2. The methods are terminated when ${\rm RES}\leq 10^{-9}$.
4.3 GRMS vs. NSOR
For NSOR, similar to the proof of Theorem 1, the smaller value of $\rho(T_{\rm NSOR})$ is, the faster the NSOR iteration (23) will converge later on, where
$T_{\rm NSOR}=\begin{bmatrix}|1-\omega_{\rm NSOR}|&\omega_{\rm NSOR}\|A^{-1}B\|\\ \frac{\omega_{\rm NSOR}|1-\omega_{\rm NSOR}|}{\sigma_{\rm NSOR}}(\|A\|+\|B\|)&\frac{\omega_{\rm NSOR}^{2}}{\sigma_{\rm NSOR}}\|A^{-1}B\|\|B\|+\|I-\frac{\omega_{\rm NSOR}^{2}}{\sigma_{\rm NSOR}}B\|\\ \end{bmatrix}.$
For GRMS and NSOR, we have the following comparison theorem.
Theorem 4.
Let $A,B\in\mathbb{R}^{n\times n}$ with $A=M-N$ such that $M$ is nonsingular. $Q=Q_{1}-Q_{2}\in\mathbb{R}^{n\times n}$ and $Q_{1}$ is nonsingular. $\sigma_{\rm NSOR}$, $\omega_{\rm NSOR}>0$ and $\theta>0$. Let notation (25) and condition (73) hold, and $T$ is an irreducible matrix.
If
$a<|1-\omega_{\rm NSOR}|,~{}c<\omega_{\rm NSOR}\|A^{-1}B\|,~{}\alpha c+d<1$ (79a)
and
$\left[\frac{\omega_{\rm NSOR}|1-\omega_{\rm NSOR}|}{\sigma_{\rm NSOR}}\left(\|A\|+\|B\|\right)\right](\alpha c+d-1)-\beta\big{[}(\alpha+1)(d-1)+\alpha c\big{]}$ (79b)
$<a\alpha\left[\frac{\omega^{2}_{\rm NSOR}}{\sigma_{\rm NSOR}}\|A^{-1}B\|\|B\|+\|I-\frac{\omega^{2}_{\rm NSOR}}{\sigma_{\rm NSOR}}B\|-1\right]$ (79c)
such that
$\hat{\mu}<\theta<\min\left\{\hat{\nu},~{}\Delta\right\},$ (79d)
where
$\hat{\mu}=\frac{\frac{\omega^{2}_{\rm NSOR}}{\sigma_{\rm NSOR}}\|A^{-1}B\|\|B\|+\|I-\frac{\omega^{2}_{\rm NSOR}}{\sigma_{\rm NSOR}}B\|-\beta c-1}{\alpha c+d-1},$
$\hat{\nu}=\frac{\frac{\omega^{2}_{\rm NSOR}}{\sigma_{\rm NSOR}}\|A^{-1}B\|\|B\|+\|I-\frac{\omega^{2}_{\rm NSOR}}{\sigma_{\rm NSOR}}B\|-\beta c+1}{1+\alpha c+d}$
and
$\Delta=\frac{\frac{\omega_{\rm NSOR}|1-\omega_{\rm NSOR}|}{\sigma_{\rm NSOR}}\big{(}\|A\|+\|B\|\big{)}-\big{(}a+1\big{)}\beta}{a\alpha},$
then $\rho(T)<\rho(T_{\rm NSOR})<1$.
Proof.
According to Lemma 2, condition (79) and the assumption that (73) hold, we know that $0\leq T<T_{\rm NSOR}$, which implied that $\rho(T)<\rho(T_{\rm NSOR})<1$ since $T$ is an irreducible matrix. ∎
The Example 3 can be used to clarify Theorem 4.
Example 3.
Consider AVEs (2) with $A=M-N$ and $b=Ax^{*}-|x^{*}|$ where $x^{*}=[\frac{1}{2},0,\frac{1}{2},0]^{\top}$.
  1. [(i)]
  2. 1.
    For GRMS iteration: $M=\begin{bmatrix}4&-0.5&0.5&0\\ -0.5&4&0&0.5\\ 0.5&0&4&-0.5\\ 0&0.5&-0.5&4\\ \end{bmatrix}$, $N=\dfrac{1}{10}I$, $Q=0.99I$, $Q_{1}=0.98I$, $H=0.01I$ and $\theta=1$. Then $T\approx\begin{bmatrix}0.0333&0.3300\\ 0.0446&0.3503\end{bmatrix}$ is an irreducible matrix and $\rho(T)\approx 0.3914$;
  3. 2.
    For NSOR iteration: $\omega_{\rm NSOR}=1.1,~{}\sigma_{\rm NSOR}=0.98$. Then $T_{\rm NSOR}\approx\begin{bmatrix}0.1000&0.3793\\ 0.6622&0.6605\end{bmatrix}$ and $\rho(T_{\rm NSOR})\approx 0.9544$.
We can find that the conditions of Theorem 4 hold, and $\rho(T)<\rho(T_{\rm NSOR})<1$. Set initial vector $x^{(0)}=y^{(0)}=[0,0,0,0]^{\top}$. From Figure 3, we know that both the GRMS method and the NSOR method are convergent and the former is perform well than the latter in the cases of IT.
Figure 3: The curve of RES:=$\frac{\|Ax^{(k)}-|x^{(k)}|-b\|_{2}}{\|b\|_{2}}$ and IT (the number of iteration) between GRMS and NSOR iteration. The experiment is terminated if RES $\leq 10^{-9}$.